# Deep Unfolding-Aided Parameter Tuning for Plug-and-Play Based Video Snapshot Compressive Imaging

### v2

@@ III-D Other Related Work
Several PnP-based approach have also been proposed for video SCI. In [29], the parameter in the denoising network is updated adaptively in the reconstruction to mitigate the difference between the training data and the reconstructed videos. Since the noise level parameters are still treated as the hyperparameters, the focused point is different from our approach for the training of the noise level parameters. We might be able to combine our proposed approach with the method in [29]. In [30], the denoising network with the structure of auto-encoder, DnCNN, U-net, or ResNet is considered. The parameters in the network are then trained via deep unfolding. In this approach, however, we have to train many parameters compared with our approach in Section IV. Moreover, we need to design much complicated denoising network to use the information between multiple frames in the video. A GAP-based network [31] achieves better reconstruction accuracy than various methods, whereas large memory consumption might be a problem for high compression rate. In [32], a reconstruction method based on deep equilibrium models has been proposed. This memory effcient approach is one of the solutions for the training of the PnP-based method. The trained parameters in this approach is unchanged over the algorithm iterations, whereas our focus is on the design of parameter values at each iteration.
@@ IV-A Proposed Network via Deep Unfolding
In this paper, we use the deep unfolding approach for the tuning of parameters $\left\{\sigma_{k}\right\}_{k=0,\ldots,K-1}$ in PnP-ADMM and PnP-GAP. A reasonable parameter values for iterative algorithms should minimize a loss function $\mathcal{L}\left(\hat{\bm{x}},\bm{x}^{\ast}\right)$, which represents the difference between the estimate $\hat{\bm{x}}$ and its true value $\bm{x}^{\ast}$. For the deep unfolding-based parameter tuning, we prepare the data $\left(\bm{y}_{(1)},\bm{x}_{(1)}^{\ast}\right),\dotsc,\left(\bm{y}_{(J)},\bm{x}_{(J)}^{\ast}\right)$ for $\left(\bm{y},\bm{x}^{\ast}\right)$ and train the parameters $\left\{\sigma_{k}\right\}_{k=0,\dotsc,K-1}$ by minimizing the loss function, where $J$ is the number of data. Specifically, we consider the optimization problem
@@
and tune the parameter on the basis of the optimization problem in (31).
For the training of the parameters $\left\{\sigma_{k}\right\}_{k=0,\dotsc,K-1}$, we need to compute the gradient of the loss function with respect to the parameters. The idea of deep unfolding is to regard the iterative algorithm as a deep neural network and to compute the gradient by backpropagation. For example, the signal flow of PnP-ADMM and its unfolded version are shown in Fig. 2. Note that some signal flows are omitted for simplicity. If the denoiser $\mathcal{D}_{\sigma_{k}}(\cdot)$ is differentiable with respect to $\sigma_{k}$, the gradient of the loss function can also be computed by backpropagation because each iteration of the algorithm can be differentiable.
@@ IV-B1 Dataset
As the training dataset, we use a modified version of the DAVIS dataset [53]. The DAVIS dataset was converted to grayscale videos and downsampled so that the height is 256 pixels. Then, the width was also cut to 256 pixels by taking the central part. For all videos, $N_{t}=8$ frames of the video were compressed into a single image. The maximum pixel value of the images is normalized to one. For each observed image, we have added additive Gaussian noise with zero mean and standard deviation of $0.01$. For the training data, we use 671 observed images (corresponding to 5,368 unknown frames) in the modified DAVIS dataset. As the test dataset, we use six grayscale videos (kobe, traffic, runner, drop, crash, aerial[21], which are usually used for the performance evaluation of video SCI algorithms. For the mask $\bm{M}$, we use the same mask as the one used in the dataset of the six grayscale videos.
@@ IV-B2 Reparametrization
where $p_{k}$ is the actual trained parameter. Note that $\sigma_{k}$ in (32) is in the range $(0,1)$. For comparison, the number of iterations in PnP-ADMM and PnP-GAP is set to 60. The initial value of $p_{k}$ in the training is set so that the noise level parameter $\sigma_{k}$ follows the first setting in Section III-C. This means
@@ IV-B3 Training
As the optimizer, we use the Adam optimizer [59] with a learning rate of $0.01$. To reduce the complexity, we do not use incremental training [51], which is often used to prevent vanishing gradient problem in deep unfolding. The number of iterations is set to $K=60$, the parameter of ADMM is $\rho=0.01$, the minibatch size is $5$, and the number of epochs is $10$. Loss function $\mathcal{L}$ is the mean squared error (MSE) between the estimated video $\bm{x}^{K}$ and the ground-truth video $\bm{x}^{\ast}$.
In the experiments, we use FastDVDnet [44] as the denoising method $\mathcal{D}_{\sigma_{k}}(\cdot)$. Since there are 60 FastDVDnet layers in the unfolded networks, we require huge GPU memory for the usual training. We thus use the checkpointing approach [48, 49] to reduce GPU memory usage, which can be implemented with a function of PyTorch. In regular backpropagation calculations, all intermediate results from the forward pass are stored for use in the backward pass. Checkpointing reduces memory usage by storing only selected intermediate results from the forward pass. For the non-stored results, recalculation is performed during the backward pass using the stored intermediate results. Consequently, this approach needs longer training time than the normal approach.
@@ V Simulation Results and Discussion
In this section, we evaluate the performance of the PnP-based methods with the proposed noise level parameter tuning via computer simulations. The aim of the simulation is to demonstrate the effect of the parameter settings, and hence we do not focus on the performance comparison with other SCI reconstruction methods. Note again that the proposed approach can be combined with some other approaches as discussed in Section III-D. The simulation is conducted by using a computer with Intel Core i9-13900, 64GB memory, and NVIDIA RTX A4500. The simulation code is implemented in PyTorch 2.2.
@@ V-A Trained Noise Level Parameters and Average Reconstruction Accuracy
In the figure, ‘proposed’ represents the noise level parameters obtained by the proposed method. For comparison, we also show the noise level parameters in the first, second, and third setting in Section III-C as ‘conventional (1)’, ‘conventional (2)’, and ‘conventional (3)’, respectively. Note that the noise level parameters in the first setting are the initial values in the training. In the simulations, we use the mask $\bm{M}$ in the dataset of the six grayscale videos. From Fig. 3, we can see that the noise level parameters change significantly by the training. The trained parameters is not necessarily monotonically decreasing, though the convergence of PnP-ADMM is usually guaranteed for a decreasing sequence of noise level parameters [50, 28, 34, 35]. As shown in later, however, the proposed noise level parameters can achieve higher reconstruction accuracy than the conventional ones. This result implies that monotonically decreasing noise level parameters might not be necessarily optimal in terms of the final reconstruction accuracy. A similar phenomena to the zigzag behavior of the parameters has also been observed in the training of the TISTA [51], which is a deep unfolding-based method for sparse signal recovery. Moreover, in the reinforcement learning-based parameter tuning for PnP methods [60], the resultant noise level parameters are not necessarily monotonically decreasing for medical resonance imaging (MRI).
@@
In the figure, ‘proposed’ represents the peak signal-to-noise ratio (PSNR) of PnP-ADMM with the noise level parameter obtained by the proposed deep unfolding approach. On the other hand, ‘conventional (1)’, ‘conventional (2)’, and ‘conventional (3)’ represent the performance of PnP-ADMM with the noise level parameters in the first, second, and third setting in Section III-C, respectively. The PSNR is calculated by averaging the result over the test data composed of six videos. The standard deviation of the Gaussian measurement noise is $0.01$, which is the same in the training. From Fig. 4, we observe that the noise level parameters significantly affect the reconstruction accuracy. We can also see that the proposed parameter tuning can achieve higher reconstruction accuracy than the conventional ones, though the PSNR is not monotonically increasing because of the zig-zag behavior of the noise level parameters.
Next, we evaluate the effectiveness of the proposed approch for PnP-GAP. Figure 5 shows the noise level parameters $\left\{\sigma_{k}\right\}_{k=0,\dotsc,K-1}$ obtained by the proposed method via deep unfolding for PnP-GAP.
@@
As in the case of PnP-ADMM, the noise level parameters change significantly during the training. The resulant noise level parameters are not necessarily monotonically decreasing and the zig-zag behavior is observed. This result suggests that the non-decreasing noise level parameters are better in terms of the final reconstruction accuracy also in the case of PnP-GAP.
Figure 6 shows the reconstruction accuracy of PnP-GAP for various parameter settings.
@@
As in the case with PnP-ADMM, the PSNR is calculated by averaging the result over the test data of six grayscale videos. The standard deviation of the Gaussian measurement noise is $0.01$. Figure 6 shows that the proposed parameter tuning can also achieve higher reconstruction accuracy than the conventional ones for PnP-GAP, though the PSNR is not monotonically increasing.
@@ V-B Reconstruction Results for Each Video
The standard deviation of the Gaussian measurement noise is $0.01$, which is the same in the training. In the figure, the top row shows the ground-truth, the middle row shows the reconstruction results of conventional first parameter setting in Section III-C, and the bottom row shows the reconstruction results of the proposed parameters. The numbers above the images are the indices of the frames. We can see that the proposed parameter tuning can achieve better reconstruction results than the conventional ones. In particular, although the conventional parameters fail to reconstruct the details of the cars, the proposed parameters can reconstruct the details more clearly.
A possible reason for the performance degradation of the conventional parameters is that the initial value of the algorithm is simply set as $\bm{v}^{0}=\bm{\Phi}^{\top}\bm{y}$ for the simple comparison. If we use the initizalization with GAP-total variation (GAP-TV) as in [28], the performance of the conventional parameters would be improved. By the proposed approch, however, we can obtain good reconstruction results without such initialization and reduce the computational cost. It should be noted that the proposed parameter tuning can also be combined with such initialization method.
TABLE I shows the PSNR and SSIM of PnP-ADMM for kobe, traffic, runner, drop, crash, and aerial.
TABLE I: PSNR (left entry) and SSIM (right entry) of PnP-ADMM. (noisy case, standard deviation: $0.01$)
@@
conventional (1) 29.83/0.924 22.73/0.857 34.24/0.946 41.02/0.981 25.16/0.894 27.40/0.880 30.06/0.914
conventional (2) 25.11/0.853 19.69/0.731 31.71/0.920 39.24/0.971 24.36/0.867 25.83/0.849 27.66/0.865
conventional (3) 22.40/0.790 17.60/0.570 24.26/0.849 16.90/0.283 17.63/0.474 20.50/0.693 19.88/0.610
@@
TABLE II shows those of PnP-GAP.
TABLE II: PSNR (left entry) and SSIM (right entry) of PnP-GAP. (noisy case, standard deviation: $0.01$)
@@
conventional (1) 25.88/0.608 21.03/0.619 25.75/0.511 26.20/0.472 23.00/0.513 23.77/0.491 24.27/0.536
conventional (2) 25.29/0.647 20.20/0.614 27.03/0.572 27.58/0.538 23.35/0.557 24.22/0.540 24.61/0.578
conventional (3) 19.42/0.457 15.37/0.365 20.40/0.444 15.69/0.182 15.82/0.282 16.22/0.353 17.15/0.347
@@
The standard deviation of the Gaussian noise is $0.01$. From the tables, we can see that the proposed parameter tuning can achieve higher reconstruction accuracy than the conventional ones for each video.
Finally, we evaluate the robustness of the proposed parameter to the standard deviation of the measurement noise. TABLE III shows the PSNR and SSIM of PnP-ADMM when the standard deviation of the Gaussian measurement noise is 00, i.e., $\bm{e}=\bm{0}$.
TABLE III: PSNR (left entry) and SSIM (right entry) of PnP-ADMM. (noiseless case)
@@
conventional (1) 29.99/0.929 23.01/0.867 34.53/0.950 41.90/0.989 25.26/0.900 27.41/0.875 30.35/0.918
conventional (2) 25.35/0.859 19.75/0.742 31.91/0.924 39.35/0.976 24.45/0.873 25.91/0.857 27.79/0.872
conventional (3) 22.41/0.796 17.59/0.572 24.06/0.847 16.89/0.283 17.63/0.477 20.50/0.697 19.85/0.612
@@
TABLE IV shows those of PnP-GAP.
TABLE IV: PSNR (left entry) and SSIM (right entry) of PnP-GAP. (noiseless case)
@@
conventional (1) 26.37/0.647 21.26/0.642 26.88/0.575 27.94/0.564 23.40/0.558 24.01/0.524 24.97/0.585
conventional (2) 25.62/0.676 20.33/0.632 27.72/0.616 28.02/0.571 23.56/0.590 24.47/0.570 24.95/0.609
conventional (3) 19.49/0.490 15.36/0.370 20.43/0.488 15.69/0.187 15.83/0.293 16.25/0.379 17.18/0.368
@@
Moreover, TABLEs V and VI show the performance of PnP-ADMM and PnP-GAP when the standard deviation of the Gaussian measurement noise is $0.05$, respectively.
TABLE V: PSNR (left entry) and SSIM (right entry) of PnP-ADMM. (noisy case, standard deviation: $0.05$)
@@
conventional (1) 26.17/0.785 20.57/0.729 30.07/0.798 32.89/0.817 23.86/0.764 26.18/0.753 26.62/0.774
conventional (2) 22.14/0.644 18.74/0.600 28.25/0.779 31.76/0.809 23.48/0.739 24.68/0.717 24.84/0.715
conventional (3) 21.88/0.671 17.46/0.515 24.11/0.734 16.88/0.272 17.50/0.419 20.46/0.621 19.72/0.539
@@
TABLE VI: PSNR (left entry) and SSIM (right entry) of PnP-GAP. (noisy case, standard deviation: $0.05$)
@@
conventional (1) 23.67/0.509 19.49/0.524 23.84/0.414 24.73/0.408 21.83/0.423 22.78/0.426 22.72/0.451
conventional (2) 22.64/0.512 18.99/0.512 25.26/0.502 26.32/0.488 22.56/0.485 23.17/0.468 23.16/0.494
conventional (3) 19.10/0.390 15.27/0.333 20.08/0.351 15.60/0.169 15.70/0.253 16.08/0.305 16.97/0.300
@@
From the tables, we can see that the proposed parameter tuning can achieve higher reconstruction accuracy than the conventional ones for most videos, even though the parameter is trained by using the data with the standard deviation of $0.01$. This result suggests that the proposed parameter tuning is robust to the change of the standard deviation of the measurement noise. Alghough the performance in the noiseless case is slightly worse than that in the noisy case with $\sigma=0.01$ for some videos, this would be because the noise level parameters are trained by using the noisy data.

### v3

@@ III-D Other Related Work
Several PnP-based approaches have also been proposed for video SCI. In [29], the parameter in the denoising network is updated adaptively in the reconstruction to mitigate the difference between the training data and reconstructed videos. Since the noise level parameters are still treated as hyperparameters, the focused point is different from our approach for training the noise level parameters. We might be able to combine our proposed approach with the method in [29]. In [30], a denoising network with the structure of auto-encoder, DnCNN, U-net, or ResNet is considered. The parameters in the network are then trained via deep unfolding. In this approach, however, we have to train many parameters compared with our approach in Section IV. Moreover, we need to design a highly complicated denoising network to use the information between multiple frames in the video. A GAP-based network [31] achieves better reconstruction accuracy than other methods, whereas large memory consumption might be a problem for a high compression rate. In [32], a reconstruction method based on deep equilibrium models has been proposed. This memory efficient approach is one solution for the training of the PnP-based method. The trained parameters in this approach remains unchanged over the algorithm iterations, whereas our focus is on the design of the parameter values at each iteration.
@@ IV-A Proposed Network via Deep Unfolding
In this study, we use the deep unfolding approach to tune the parameters $\left\{\sigma_{k}\right\}_{k=0,\ldots,K-1}$ in PnP-ADMM and PnP-GAP. Reasonable parameter values for iterative algorithms should minimize a loss function $\mathcal{L}\left(\hat{\bm{x}},\bm{x}^{\ast}\right)$, which represents the difference between the estimate $\hat{\bm{x}}$ and its true value $\bm{x}^{\ast}$. For the deep unfolding-based parameter tuning, we prepare the data $\left(\bm{y}_{(1)},\bm{x}_{(1)}^{\ast}\right),\dotsc,\left(\bm{y}_{(J)},\bm{x}_{(J)}^{\ast}\right)$ for $\left(\bm{y},\bm{x}^{\ast}\right)$ and train the parameters $\left\{\sigma_{k}\right\}_{k=0,\dotsc,K-1}$ by minimizing the loss function, where $J$ is the number of data. Specifically, we consider the optimization problem
@@
and tune the parameter based on the optimization problem in (31).
To train the parameters $\left\{\sigma_{k}\right\}_{k=0,\dotsc,K-1}$, we need to compute the gradient of the loss function with respect to the parameters. The idea of deep unfolding is to regard the iterative algorithm as a deep neural network and compute the gradient by backpropagation. For example, the signal flow of PnP-ADMM and its unfolded version are shown in Fig. 2. Note that some signal flows are omitted for simplicity. If the denoiser $\mathcal{D}_{\sigma_{k}}(\cdot)$ is differentiable with respect to $\sigma_{k}$, the gradient of the loss function can also be computed by backpropagation, because each iteration of the algorithm can be differentiable.
@@ IV-B1 Dataset
As the training dataset, we use a modified version of the DAVIS dataset [53]. The DAVIS dataset was converted into grayscale videos and downsampled so that the height is 256 pixels. The width was then cut to 256 pixels by taking the central part. For all videos, $N_{t}=8$ frames of the video were compressed into a single image. The maximum pixel value of the images is normalized to one. For each observed image, we added additive Gaussian noise with zero mean and standard deviation of $0.01$. For the training data, we use 671 observed images (corresponding to 5,368 unknown frames) in the modified DAVIS dataset. As the test dataset, we use six grayscale videos (kobe, traffic, runner, drop, crash, aerial[21], which are typically used for the performance evaluation of video SCI algorithms. For the mask $\bm{M}$, we use the same mask as that in the dataset of the six grayscale videos.
@@ IV-B2 Reparametrization
where $p_{k}$ is the actual trained parameter. Note that $\sigma_{k}$ in (32) is within the range $(0,1)$. For comparison, the number of iterations in PnP-ADMM and PnP-GAP is set to 60. The initial value of $p_{k}$ in the training is set such that the noise level parameter $\sigma_{k}$ follows the first setting in Section III-C. This means
@@ IV-B3 Training
As the optimizer, we use the Adam optimizer [59] with a learning rate of $0.01$. To reduce computational complexity, we do not use incremental training [51], which is often used to prevent the vanishing gradient problem in deep unfolding. The number of iterations is set to $K=60$, the parameter of ADMM is $\rho=0.01$, the minibatch size is $5$, and the number of epochs is $10$. The loss function $\mathcal{L}$ is the mean squared error (MSE) between the estimated video $\bm{x}^{K}$ and the ground-truth video $\bm{x}^{\ast}$.
In the experiments, we use FastDVDnet [44] as the denoising method $\mathcal{D}_{\sigma_{k}}(\cdot)$. Since there are 60 FastDVDnet layers in the unfolded networks, we require huge GPU memory for the usual training. We thus use the checkpointing approach [48, 49] to reduce GPU memory usage, which can be implemented with a function of PyTorch. In regular backpropagation calculations, all the intermediate results from the forward pass are stored for use in the backward pass. Checkpointing reduces memory usage by storing only selected intermediate results from the forward pass. For the non-stored results, recalculation is performed during the backward pass using the stored intermediate results. Consequently, this approach requires longer training time than the normal approach.
@@ V Simulation Results and Discussion
In this section, we evaluate the performance of the PnP-based methods with the proposed noise level parameter tuning via computer simulations. Although there are various methods for the SCI reconstruction, the aim of the simulation here is to demonstrate the effect of the parameter settings. Therefore, we do not focus on a performance comparison with other SCI reconstruction methods. Note again that the proposed approach can be combined with some other approaches, as discussed in Section III-D. The simulation is conducted by using a computer with Intel Core i9-13900, 64GB memory, and NVIDIA RTX A4500. The simulation code is implemented in PyTorch 2.2.
@@ V-A Trained Noise Level Parameters and Average Reconstruction Accuracy
In the figure, ‘proposed’ represents the noise level parameters obtained by the proposed method. For comparison, we also show the noise level parameters in the first, second, and third setting in Section III-C as ‘conventional (step)’, ‘conventional (exponential)’, and ‘conventional (constant)’, respectively. Note that the noise level parameters in the first setting are the initial values in the training. In the simulations, we use the mask $\bm{M}$ in the dataset of the six grayscale videos. From Fig. 3, we can see that the noise level parameters change significantly by the training.
The trained parameters is not necessarily monotonically decreasing, though the convergence of PnP-ADMM is usually guaranteed for a decreasing sequence of noise level parameters [50, 28, 34, 35]. As shown in later, however, the proposed noise level parameters can achieve a higher reconstruction accuracy than the conventional ones. This result implies that monotonically decreasing noise level parameters might not necessarily be optimal in terms of the final reconstruction accuracy. A phenomenon similar to the zigzag behavior of the parameters has also been observed in the training of the TISTA [51], which is a deep unfolding-based method for sparse signal recovery. Moreover, in reinforcement learning-based parameter tuning for PnP methods [60], the resultant noise level parameters are not necessarily monotonically decreasing for medical resonance imaging (MRI).
@@
In the figure, ‘proposed’ represents the peak signal-to-noise ratio (PSNR) of PnP-ADMM with the noise level parameter obtained by the proposed deep unfolding approach. On the other hand, ‘conventional (step)’, ‘conventional (exponential)’, and ‘conventional (constant)’ represent the performance of PnP-ADMM with the noise level parameters in the first, second, and third setting in Section III-C, respectively. The PSNR is calculated by averaging the result over the test data composed of six videos. For all videos, $N_{t}=8$ frames of the video are reconstructed from a single measurement image. The standard deviation of the Gaussian measurement noise is $0.01$, which is the same in the training. From Fig. 4, we observe that the noise level parameters significantly affect the reconstruction accuracy. We can also see that the proposed parameter tuning can achieve higher reconstruction accuracy than the conventional ones, though the PSNR is not monotonically increasing because of the zig-zag behavior of the noise level parameters.
Next, we evaluate the effectiveness of the proposed approach for PnP-GAP. Figure 5 shows the noise level parameters $\left\{\sigma_{k}\right\}_{k=0,\dotsc,K-1}$ obtained by the proposed method via deep unfolding for PnP-GAP.
@@
As in the case of PnP-ADMM, the noise level parameters change significantly by the training. The resultant noise level parameters are not necessarily monotonically decreasing and a zig-zag behavior is observed. This result suggests that the non-decreasing noise level parameters are better in terms of the final reconstruction accuracy also in the case of PnP-GAP.
Figure 6 shows the reconstruction accuracy of PnP-GAP for various parameter settings. As in the case with PnP-ADMM, the PSNR is calculated by averaging the results over the test data of six grayscale videos. The standard deviation of the Gaussian measurement noise is $0.01$. Figure 6 shows that the proposed parameter tuning can also achieve a higher reconstruction accuracy than the conventional ones for PnP-GAP, though the PSNR is not increasing monotonically.
@@ V-B Reconstruction Results for Each Video
In Figs. 7 and 8, we show the reconstruction progress of PnP-ADMM per iteration for a frame of kobe and crash, respectively.
Figure 7: Reconstruction progress of PnP-ADMM per iteration for a frame of kobe.
(PSNR (dB)/SSIM of conventional (step): 25.81/0.884, PSNR (dB)/SSIM of proposed: 29.89/0.908)
Figure 8: Reconstruction progress of PnP-ADMM per iteration for a frame of crash.
(PSNR (dB)/SSIM of conventional (step): 23.39/0.857, PSNR (dB)/SSIM of proposed: 26.78/0.905)
The standard deviation of the Gaussian measurement noise is $0.01$. In the figures, the top row shows the estimate at the $k$-th iteration ($k=10,20,\dotsc,60$) with the conventional first parameter setting (step) and the corresponding ground truth image. The bottom row shows the estimate obtained using the proposed noise level parameters. From the figures, the reconstructed result of PnP-ADMM improves as the iteration proceeds. We also observe that the proposed noise level parameters can reconstruct the details of the scene better than the conventional ones. For the frame of kobe, the PSNR and structural similarity index measure (SSIM) at $k=60$ for the conventional parameters are 25.81 dB and 0.884, respectively, whereas they are 29.89 dB and 0.908 for the proposed parameters. For the frame of crash, the PSNR and SSIM at $k=60$ for the conventional parameters are 23.39 dB and 0.857, respectively, whereas 26.78 dB and 0.905 for the proposed parameters. From the quantitative evaluation, we can also see that the proposed noise level parameters can achieve better reconstruction accuracy than the conventional ones.
@@
The standard deviation of the Gaussian measurement noise is $0.01$, which is the same as that in the training. In the figure, the top row shows the ground-truth, the middle row shows the reconstruction results of the conventional first parameter setting (step) in Section III-C, and the bottom row shows the reconstruction results of the proposed parameters. The numbers above the images represent the indices of the frames. We can see that the proposed parameter tuning can achieve better reconstruction results than conventional ones. In particular, although the conventional parameters fail to reconstruct the details of the cars, the proposed parameters can reconstruct the details more clearly.
A possible reason for the performance degradation of the conventional parameters is that the initial value of the algorithm is simply set as $\bm{v}^{0}=\bm{\Phi}^{\top}\bm{y}$ for the simple comparison. If we use the initialization with GAP-total variation (GAP-TV) as in [28], the performance of the conventional parameters would be improved. By the proposed approach, however, we can obtain good reconstruction results without such initialization and reduce the computational cost. It should be noted that the proposed parameter tuning can also be combined with such initialization method.
TABLE II shows the PSNR and SSIM of PnP-ADMM for kobe, traffic, runner, drop, crash, and aerial.
TABLE II: PSNR (left entry) and SSIM (right entry) of PnP-ADMM. (noisy case, standard deviation: $0.01$)
@@
conventional (step) 29.83/0.924 22.73/0.857 34.24/0.946 41.02/0.981 25.16/0.894 27.40/0.880 30.06/0.914
conventional (exponential) 25.11/0.853 19.69/0.731 31.71/0.920 39.24/0.971 24.36/0.867 25.83/0.849 27.66/0.865
conventional (constant) 22.40/0.790 17.60/0.570 24.26/0.849 16.90/0.283 17.63/0.474 20.50/0.693 19.88/0.610
@@
TABLE III lists the values of PnP-GAP.
TABLE III: PSNR (left entry) and SSIM (right entry) of PnP-GAP. (noisy case, standard deviation: $0.01$)
@@
conventional (step) 25.88/0.608 21.03/0.619 25.75/0.511 26.20/0.472 23.00/0.513 23.77/0.491 24.27/0.536
conventional (exponential) 25.29/0.647 20.20/0.614 27.03/0.572 27.58/0.538 23.35/0.557 24.22/0.540 24.61/0.578
conventional (constant) 19.42/0.457 15.37/0.365 20.40/0.444 15.69/0.182 15.82/0.282 16.22/0.353 17.15/0.347
@@
The standard deviation of the Gaussian noise is $0.01$. From the tables, we can see that the proposed parameter tuning can achieve a higher reconstruction accuracy than the conventional parameters for each video.
Finally, we evaluate the robustness of the proposed parameter against the standard deviation of the measurement noise. TABLE IV shows the PSNR and SSIM of PnP-ADMM when the standard deviation of the Gaussian measurement noise is 00, that is, $\bm{e}=\bm{0}$.
TABLE IV: PSNR (left entry) and SSIM (right entry) of PnP-ADMM. (noiseless case)
@@
conventional (step) 29.99/0.929 23.01/0.867 34.53/0.950 41.90/0.989 25.26/0.900 27.41/0.875 30.35/0.918
conventional (exponential) 25.35/0.859 19.75/0.742 31.91/0.924 39.35/0.976 24.45/0.873 25.91/0.857 27.79/0.872
conventional (constant) 22.41/0.796 17.59/0.572 24.06/0.847 16.89/0.283 17.63/0.477 20.50/0.697 19.85/0.612
@@
TABLE V shows those of PnP-GAP.
TABLE V: PSNR (left entry) and SSIM (right entry) of PnP-GAP. (noiseless case)
@@
conventional (step) 26.37/0.647 21.26/0.642 26.88/0.575 27.94/0.564 23.40/0.558 24.01/0.524 24.97/0.585
conventional (exponential) 25.62/0.676 20.33/0.632 27.72/0.616 28.02/0.571 23.56/0.590 24.47/0.570 24.95/0.609
conventional (constant) 19.49/0.490 15.36/0.370 20.43/0.488 15.69/0.187 15.83/0.293 16.25/0.379 17.18/0.368
@@
Moreover, TABLEs VI and VII show the performance of PnP-ADMM and PnP-GAP when the standard deviation of the Gaussian measurement noise is $0.05$, respectively.
TABLE VI: PSNR (left entry) and SSIM (right entry) of PnP-ADMM. (noisy case, standard deviation: $0.05$)
@@
conventional (step) 26.17/0.785 20.57/0.729 30.07/0.798 32.89/0.817 23.86/0.764 26.18/0.753 26.62/0.774
conventional (exponential) 22.14/0.644 18.74/0.600 28.25/0.779 31.76/0.809 23.48/0.739 24.68/0.717 24.84/0.715
conventional (constant) 21.88/0.671 17.46/0.515 24.11/0.734 16.88/0.272 17.50/0.419 20.46/0.621 19.72/0.539
@@
TABLE VII: PSNR (left entry) and SSIM (right entry) of PnP-GAP. (noisy case, standard deviation: $0.05$)
@@
conventional (step) 23.67/0.509 19.49/0.524 23.84/0.414 24.73/0.408 21.83/0.423 22.78/0.426 22.72/0.451
conventional (exponential) 22.64/0.512 18.99/0.512 25.26/0.502 26.32/0.488 22.56/0.485 23.17/0.468 23.16/0.494
conventional (constant) 19.10/0.390 15.27/0.333 20.08/0.351 15.60/0.169 15.70/0.253 16.08/0.305 16.97/0.300
@@
From the tables, we can see that the proposed parameter tuning can achieve higher reconstruction accuracy than the conventional ones for most videos, even though the parameter is trained by using the data with the standard deviation of $0.01$. This result suggests that the proposed parameter tuning is robust to changes in the standard deviation of the measurement noise. Although the performance in the noiseless case is slightly worse than that in the noisy case with $\sigma=0.01$ for some videos, this would be because the noise level parameters are trained by using the noisy data.